# LLM Agents for Interactive Workflow Provenance: Reference Architecture and Evaluation Methodology

### v1

@@ Abstract.
This is a preprint of a paper accepted for publication in the proceedings of the Supercomputing Conference. Please cite using the ACM Reference Format above.
@@ 1. Introduction
This paper introduces a modular, loosely coupled provenance agent system architecture that enables live interaction between users and data during workflow execution in the ECH continuum. The architecture leverages LLM-powered agents and the Model Context Protocol (MCP) (mcp, 2025) to support natural language interactivity. It adopts a lightweight, metadata-driven design that translates natural language into accurate runtime workflow provenance queries. Specifically, we make the following contributions:
@@ 2.3. Reference Architecture for Distributed Workflow Provenance
The architecture prioritizes interoperability and deployment flexibility (Souza et al., 2023) through a modular, adapter-based design that supports provenance capture via two complementary mechanisms: (i) non-intrusive observability adapters, which passively monitor dataflow from services such as RabbitMQ, SQLite, MLflow, and file systems without modifying application code; and (ii) direct code instrumentation, which uses lightweight hooks such as Python decorators to capture fine-grained task-level metadata from functions, scripts, or tools like Dask, PyTorch, and agents behind MCP. These mechanisms provide flexible entry points for capturing provenance across heterogeneous workflows. To reduce interference with HPC applications, provenance messages are buffered in-memory and streamed asynchronously to a central hub using a publish-subscribe protocol with configurable flushing strategies. Listing 1 shows an example of such a message.
@@
Code Listing 1: Example of a workflow task provenance message from the chemistry workflow used in this work.
@@ 3. LLM-Powered Provenance Agent Evaluation Methodology
An alternative is to follow lightweight approaches based on schema or metadata instead of full data, while partial data are still possible. Practical strategies to augment the context are: common workflow schema, which adds descriptions for domain-agnostic common task fields (e.g., task_id, activity_id); application-specific dataflow fields, which include domain-specific parameters and outputs; and semantic descriptions and domain-specific values, which expand the context with field descriptions and representative values (which refers to the partial data strategy). For instance, in climate simulation workflows for modeling Earth’s surface weather, adding sample values like 0–110 for a field named “temperature” can help the LLM infer both the range of plausible values and the likely unit (e.g., Fahrenheit), even if the user does not specify it explicitly.
@@ 4.2. Architecture Overview
Figure 4. Provenance AI Agent Design.
@@
  • In-memory context, a buffer of recent workflow task provenance messages.
@@ 5.2. LLM Agent Response Analysis
Figure 8. Impact of contextual information components against performance and token consumption.
@@
Figure 9. Impact of contextual information components against performance and token Consumption.
@@ 5.4. Evaluation Summary, Lessons Learned, and Future Work
The dynamic dataflow schema was key in this generalization, enabling the agent to track workflow structure and key parameter semantics incrementally, while avoiding context window overflows even in an HPC environment. Because the agent operates on metadata rather than provenance task data, its LLM performance is independent of the number of workflow tasks, volume of processed data, or volume of captured provenance, depending instead on the workflow complexity alone, i.e., the number and diversity of activities and their input and output fields. This trade-off allows for lightweight operation in large workflows. These results highlight the approach’s potential as a low-barrier, interactive tool to explore and visualize provenance data without requiring structured query languages, custom code, or dealing with the complexities of setting up their own data analysis environments. We believe that this work has the potential to accelerate scientific discovery by reducing the time-to-insights in complex ECH workflows.
@@
Lessons from Prompting and Evaluation. We assessed how each contextual component (i.e., query guidelines, prompt elements, dataflow schema, and domain values) contributed to individual query classes, understanding which component we should invest further in to improve the responses’ quality. OLAP queries proved the hardest across all LLMs. The query guidelines and few-shot examples helped reduce syntax errors, while the dataflow schema and domain-specific examples enabled better semantic alignment.
Our goal was not to benchmark specific LLMs, but to evaluate a metadata, query-driven approach that remains robust as models change. Across the tested models, the same pattern holds: query guidelines, the dynamic dataflow schema, and few-shot examples drive most gains, while having metadata as input and queries as output keep token usage bounded. This design should maintain high scores as newer LLMs appear, without altering our proposed architecture and approach. Even as LLMs improve and context windows grow with newer LLMs, their invocations will not ingest a whole large provenance database; Thus, our approach will remain necessary for future LLM versions, scaling by conditioning on schema and queries instead of prompting the entire database.
@@ 6. Conclusions
Our work demonstrates that LLM-powered agents, when guided by structured schema representations, dynamic prompting strategies, and a modular system architecture, can enable effective, real-time interaction with complex workflow provenance data across the ECH continuum. By decoupling LLM interaction from raw data volume and focusing on metadata, our approach remains lightweight and scalable, even for large HPC workflows. The agent generalized well from synthetic to real-world settings, achieving high accuracy without domain-specific tuning. These results suggest that interactive, provenance-aware AI agents can significantly reduce the effort required for exploratory data analysis, anomaly diagnosis, and monitoring, closing the gap between scientists and their data. We believe this work lays the foundation for a new class of intelligent provenance agents while highlighting open challenges such as dynamic semantic enrichment of schemas, feedback-driven prompt tuning, and scalable graph-based provenance querying.
Acknowledgments. ChatGPT-4o was used to help polish writing, improve conciseness, and check grammar across the sections of the paper. This research used resources of the Oak Ridge Leadership Computing Facility at Oak Ridge National Laboratory, supported by the U.S. Department of Energy Office of Science under Contract No. DE-AC05-00OR22725. It was also supported in part by an appointment to the Education Collaboration at ORNL (ECO) Program, sponsored by the DOE and administered by the Oak Ridge Institute for Science and Education, and the U.S. Department of Energy Office of Science under Contract No. DE-AC02-06CH11357

### v2

@@ Abstract.
This is a preprint for a publication in the proceedings of the Supercomputing Conference. Please cite using the ACM Reference Format (on the top-right).
@@ 1. Introduction
This paper introduces a modular, loosely coupled provenance agent system architecture that facilitates live interaction between users and data during workflow execution in the ECH continuum. The architecture leverages LLM-powered agents and the Model Context Protocol (MCP) (mcp, 2025) to support natural language interactivity. It adopts a lightweight, metadata-driven design that translates natural language into accurate runtime workflow provenance queries. Specifically, we make the following contributions:
@@ 2.3. Reference Architecture for Distributed Workflow Provenance
The architecture prioritizes interoperability and deployment flexibility (Souza et al., 2023) through a modular, adapter-based design that supports provenance capture via two complementary mechanisms: (i) non-intrusive observability adapters, which passively monitor dataflow from services such as RabbitMQ, SQLite, MLflow, and file systems without modifying application code; and (ii) direct code instrumentation, which uses lightweight hooks such as Python decorators to capture fine-grained task-level metadata from functions, scripts, or packages like PyTorch and MCP SDK, providing flexible entry points for capturing provenance across heterogeneous workflows. To reduce interference with HPC applications, provenance messages are buffered in-memory and streamed asynchronously to a central hub using a publish-subscribe protocol with configurable flushing strategies. Listing 1 shows an example of such a message.
@@
Listing 1: Example of a workflow task provenance message from the chemistry workflow used in this work.
@@ 3. LLM-Powered Provenance Agent Evaluation Methodology
An alternative is to follow lightweight approaches based on schema or metadata instead of full data. Practical strategies to augment the context are: common workflow schema, which adds descriptions for domain-agnostic common fields (e.g., task_id, activity_id); application-specific dataflow fields, which include domain-specific parameters and outputs; and semantic descriptions and domain-specific values, which expand the context with field descriptions and representative example values (which refers to the partial data strategy). For instance, in Earth’s climate workflows, adding sample values like 0–110 for a field named “temperature” can help inferring range of plausible values or the likely unit (e.g., Fahrenheit), even if the user does not specify it.
@@ 4.2. Architecture Overview
Figure 4. Provenance AI agent architecture design.
@@
  • In-memory context, which is a buffer of recent workflow task provenance messages.
@@ 5.2. LLM Agent Response Analysis
Figure 8. Impact of contextual components on performance and token consumption. See Table 2 for labels.
@@
Figure 9. Impact of contextual components on performance and data types. See Table 2 for labels.
@@ 5.4. Evaluation Summary, Lessons Learned, and Future Work
The dynamic dataflow schema was key in this generalization, enabling the agent to track workflow structure and key parameter semantics incrementally, while avoiding context window overflows even in an HPC environment. Because the agent operates on metadata rather than task provenance data, its LLM performance is independent of the number of workflow tasks, volume of processed data, or volume of captured provenance, depending instead on the workflow complexity alone, i.e., the number and diversity of activities and their input and output fields. This trade-off allows for lightweight operation in large workflows. These results highlight the approach’s potential as a low-barrier, interactive tool to explore and visualize provenance data without requiring structured query languages, custom code, or dealing with the complexities of setting up their own data analysis environments. We believe that this work has the potential to accelerate scientific discovery by reducing the time-to-insights in complex ECH workflows.
@@
Lessons from Prompting and Evaluation. We assessed how each contextual component (i.e., query guidelines, dynamic dataflow schema) contributed to individual query classes, understanding which component we should invest further in to improve the responses’ quality. OLAP queries proved the hardest across all LLMs. The query guidelines and few-shot examples helped reduce syntax errors, while the dataflow schema and domain-specific examples enabled better semantic alignment.
Our goal was not to benchmark specific LLMs, but to evaluate a metadata, query-driven approach that remains robust as models change. Across the tested models, the same pattern holds: query guidelines, the dynamic dataflow schema, and domain values drive most gains, while having metadata as input and queries as output keep token usage bounded. This design should maintain high scores as newer LLMs appear, without altering our proposed architecture and approach. Even as LLMs improve and context windows grow with newer LLMs, their invocations will not ingest a whole provenance database, which may reach gigabytes or terabytes in large runs. Thus, our approach will remain necessary for future LLM versions, scaling by depending on the dataflow schema and queries instead of prompting the entire database.
@@ 6. Conclusions
Our work demonstrates that LLM-powered agents, when guided by structured schema representations, dynamic prompting strategies, and a modular system architecture, can enable effective, near real-time interaction with complex workflow provenance data across the ECH continuum. By decoupling LLM interaction from raw data volume and focusing on metadata, our approach remains lightweight and scalable, even for large HPC workflows. The agent generalized well from synthetic to real-world settings, achieving high accuracy without domain-specific tuning. These results suggest that interactive, provenance-aware AI agents can significantly reduce the effort required for exploratory data analysis, anomaly diagnosis, and monitoring, closing the gap between scientists and their data. We believe this work lays the foundation for a new class of intelligent provenance agents while highlighting open challenges such as dynamic semantic enrichment of schemas, feedback-driven autonomous prompt tuning, and graph-based LLM querying.
Acknowledgments. ChatGPT-4o was used to help polish writing, improve conciseness, and check grammar across the sections of the paper. This research used resources of the Oak Ridge Leadership Computing Facility at Oak Ridge National Laboratory, supported by the U.S. Department of Energy Office of Science under Contract No. DE-AC05-00OR22725. It was also supported in part by an appointment to the Education Collaboration at ORNL (ECO) Program, sponsored by the DOE and administered by the Oak Ridge Institute for Science and Education, and the U.S. Department of Energy Office of Science under Contract No. DE-AC02-06CH11357.